OpenStudy (anonymous):

Okay am I on the right path? -15t+4=-26 -15t - 4=-26-4 -19t=-30 t= To me it doesn't seem right... I suck at Algebra.

OpenStudy (anonymous):

-15t+4=-26 collect like terms -15t=-26-4 simplify -15t=-30 divide by -15 to find t t=2

OpenStudy (anonymous):

okay so I don't subtract 4 from -15?

OpenStudy (texaschic101):

no...because they are not like terms....4 and -15t are not like terms..it would have to be 4t for it to be like terms where you would subtract

OpenStudy (anonymous):

okay thank you!

OpenStudy (texaschic101):

sure thing :)